Determine the molality of each of the following solutions




a. 2.34 mol KCI in 2.300 L of solution



b. 0.56 H2SO4 in 2 L of solution



c. 20.45 g AI (NO3)3 in 2000 mL of solution

1
Expert's answer
2022-04-08T08:44:02-0400

Molality is a concentration in mole/1000g.

Let dencity of solutions eaual 1 g/ml

a.c=2.34 /2,3=1.02 mol/kg

b. Let it be 0.56 mol ( it is not mentioned in task) of "H_2SO_4" .

c=0.56/2=0.28 mol/kg

c."M(Al(NO_3)_3)=27+3(14+16x3)=213 g\/mol"

"n(Al(NO_3)_3)=20.45\/213=0.096 mol"

"c(Al(NO_3)_3)=0.096\/2=0.048 mol\/kg"